AM radio station broadcasts at 820 kHz while its FM partner broadcasts at 89.7 MHz

Calculate the smallest increment of energy that can be emitted or absorbed at these wavelengths.


1
Expert's answer
2021-03-29T06:09:03-0400

Solution:

"h1 = \frac{c}{ν}= \frac{3×10^8 }{8.97×10^7}= 3.34m"


"h2 = \frac{c}{ν}= \frac{3×10^8}{8.2×10^5}= 365.85 m"


Answer: 3.34m, 365.85m